# A Non-Abelian Gauge Theory for Surface Excitations of ${}^{3}$e-B

### v1

@@ Abstract
In ${}^{3}$e-B, two atoms pair in an orbital angular momentum $1$ spin triplet state above the phase transition temperature with $SO(3)\times SO(3)$ symmetry. Below the transition temperature, this symmetry is spontaneously broken to the diagonal $SO(3)$ due to spin-orbit coupling. Considerations based on effective potentials and solitons show that $SO(3)$’s gets enhanced to $SU(3)$’s and the symmetry breaking is that of $G=SU(3)\times SU(3)\times U(1)$ to $H=SU(3)$. The theory of the resultant Goldtsone modes can be naturally formulated as a gauge theory of $H$. Its Gauss law is treated here and shown to lead to surface states in a container with a dynamics governed by large gauge transformations.
Observable consequences are pointed. The transference of the analysis to the chiral model of QCD is pointed out where $SU(3)$ are the left and right chiral groups, $U(1)$ is the axial $U(1)$ , the surviving symmetry is flavour $SU(3)$ and the Goldstone modes are the pions.
1 Introduction
In ${}^{3}$e-B , the superfluid transition takes place by a pairing of this molecule in its $P$-wave spin triplet state with another such molecule in an identical state.
The order parameter for this transition can be denoted by a complex $3\times 3$ matrix of fields $\phi=\{\phi_{s}(x)\}$ following Mermin [1] where $SO(3)$ of spin acts on $\phi$ by left multiplication and $SO(3)$ of orbital momentum acts on $\phi$ right multiplication.
@@ 1 Introduction
$\lim_{|\vec{x}|\rightarrow\infty}\phi(x)=\mathbb{I}$ (1.1)
@@
This symmetry breaking pattern is reminiscent of the chiral symmetry group $SU(3)_{L}\times SU(3)_{R}$ of massless $QCD$ with three flavours and its spontaneous breakdown to the diagonal $SU(3)_{L+R}$. We can in fact bring this resemblance closer by noting that $G$-invarint potentials are traces of polynomials in $\phi^{\dagger}\phi$ of the form
$\Sigma_{n}a_{n}\mathrm{Tr}(\phi^{\dagger}\phi)^{n}.$ (1.2)
These potentials are invariant if $SO(3)_{L,R}$ is enlarged to $\mathcal{G}=SU(3)_{L}\times SU(3)_{R}\times U(1)$ where $U(1)$ acts by an overall phase: $\phi\rightarrow\omega\phi,|\omega|=1$. Then the vacuum symmetry group is $H=SU(3)_{L+R}$. These are our groups of interest.
@@
$G$ does not act effectly on $\phi$. If $c=e^{\frac{2\pi i}{3}\mathbb{I}_{3}}$, with $\mathbb{I}$ being the $3\times 3$ identity matrix of $SU(3)$, generates the center of $SU(3)$, then the order $9$ discrete group $Z_{9}$ with elements diag$\{c^{m},\bar{c}^{n},\bar{c}^{n-m}\}$ with $n,m$ belonging to the set $\{1,2,3\}$ acts trivially on $\phi$. Thus in quantum theory only representations of $\mathcal{G}$ where $Z_{9}$ becomes identity are allowed. With this proviso, we can work with $G$.
In Mermin’s review , $\phi$ is restricted to $\omega\times R$ where $R$ is real and orthogonal. But a previous work [2] has found that for discussions with solitons of ${}^{3}$e-B it is important to consider $\phi$ to be any $3\times 3$ matrix ( with the boundary condition mentioned above). This paper will proceed in this latter direction.
2 Gauge Theory of Goldstone Modes
@@
3 Spontaneous Breakdown as a Gauge Theory
@@
The gauge groups $\mathcal{G},\mathcal{H}$ are the maps from spacetime to Lie groups $G,H$ in our considerations. , they are the restrictions of $\mathcal{G},\mathcal{H}$ to constant time, say 00 , that will be relevant. They are denoted by the same symbols. We consider four dimensional $M$ so that the constant time slices $M_{3}$, are three dimensional. It can be $\mathbb{R}^{3}$ or the 3-ball $B_{3}$ with coordinates $x$ and the boundary $S^{2}$. The evaluation of their elements at $x$ will be denoted by $g(x)$ and $h(x)$, respectively. Often we will suppress $x$.
@@
$hT_{\alpha}h^{-1}=h_{\beta,\alpha}T_{\beta}$ (3.1)
@@
$v=\frac{1}{2}T(\alpha)\mathrm{Tr}(vT(\alpha)),$ (3.4)
@@
$A_{\mu}(g(x))=\frac{i}{2}T(\alpha)\mathrm{~{}Tr~{}}T(\alpha)g^{-1}(x)\partial_{\mu}g(x)$ (3.6)
@@
$A_{\mu}(gh)=h^{-1}A_{\mu}h+ih^{-1}\partial_{\mu}h.$
This follows on using the orthogonality of $Ad~{}H$ and the identity (3.1) with $v_{\mu}=h^{-1}\partial_{\mu}h$ ( $x$ has been suppressed).
In contrast
$B_{\mu}(g)=\frac{i}{2}S(j)\mathrm{~{}Tr~{}}S(j)g^{-1}(x)\partial_{\mu}g(x)$ (3.8)
@@
$B_{\mu}(gh)=h^{-1}(x)B_{\mu}(g)h(x)$ (3.9)
We can now consider $\mathcal{H}$ invariant interactions like
$\mathrm{tr}F~{}^{*}F,\qquad\mathrm{Tr~{}}B_{\mu}B^{\mu}$ (3.10)
@@
$g_{L}g^{\dagger}_{R}\equiv\phi=\left\{\phi_{s,i}\right\}$ (3.11)
@@
4 Gauss Law and Superselection Rules
@@
$S=-\frac{1}{4g^{2}}\mathrm{Tr}F_{\mu\nu}F^{\mu\nu}+q\mathrm{Tr}A^{\mu}J_{\mu}$ (4.1)
@@
In view of ( 3.9) , $F$ transforms as $h^{-1}Fh$ under an $\mathcal{H}$ gauge transformation so that $S$ is invariant even if $h$ is time dependent.
@@
The derivation of this constraint by quantising $g$ in the action is awkward. But we can get it simply by varying $g$ with an infinitesimal time dependent gauge transformation:
$\delta g=g\epsilon,~{}~{}\epsilon~{}\textrm{time dependent}$ (4.2)
@@
$\epsilon=\epsilon^{\alpha}(x)\cdot T(\alpha),$ (4.3)
@@
Only $A_{0}$ varies with $\partial_{0}\epsilon$ and that variations is
$\delta A_{0}=-(\partial_{0}\epsilon^{\alpha})T(\alpha).$ (4.4)
The corresponding variation in the Lagrangian is
$\frac{1}{2g^{2}}\int d^{3}x~{}tr(D_{i}F_{i0})(\partial_{0}\epsilon^{\alpha})T(\alpha).$ (4.5)
@@
$\int d^{3}xA^{\mu}J_{\mu}$ (4.7)
to the Lagrangian which will enhance (4.6) by the matter term $J_{0}$.
We must check that L.H.S of (4.6) generates gauge transformations. and that requires the commutators of $A^{i}$ and $F_{0i}$. We can derive them as follows.
The analogue of the $pdq$ term can be read off from (4.1) . It comes from $\frac{1}{2g^{2}}\mathrm{tr}F_{0i}\partial_{0}A_{i}$ from varying $A_{i}$ so that as in $QCD$, $A^{i}$ and $E_{i}=\frac{1}{2g^{2}}F_{0i}$ are conjugate.
@@
$Q_{H}(X)=\int d^{3}x\mathrm{Tr}(D_{i}XE_{i}+XJ_{0})$ (4.8)
[…]
@@
The classical Gauss law (4.6) has the quantum version of ‘small ’ gauge tranformations with generators $Q_{H}(X)$ which fulfill a) and b):
@@
$\left[Q_{H}(X),K\right]=0.$ (4.11)
Then, $X$ being compactly supported , we recover classical Gauss law by partial integration , while (4.11) is to ensure that observables preserve the domain defined by (4.10). ( There are technical issues with domains which remain to be answered.)
@@
$Q(\Xi)=\int d^{3}x\mathrm{Tr}(D_{i}\Xi E_{i}+\Xi J_{0})$ (4.12)
where $\Xi$ are smooth functions which may not vanish at infinity ( or the boundary). These generate ‘large gauge transformations’ . As shown in previous papers, $Q(\Xi)$ commute with local observables due to local commutation relations , but in non-abelian theories such as ${}^{3}$e $B$ or $QCD$, they may not mutually commute. This has specific consequences for surface states as we shall see.
@@
    $\left[Q(\Xi),Q(\Xi^{\prime})\right]=iQ([\Xi,\Xi^{\prime}]).$ (4.13)
@@
    Adding a small gauge transformation $Q_{H}(X)$ to $Q(\Xi)$ is still a large gauge transformation $Q(\Xi+X)$ which has the same action on a quantum state in view of (4.10):
    $Q(\Xi+X)|\cdot\rangle=Q(X)|\cdot\rangle$ (4.14)
    Further since $[\Xi,X]$ is a compactly supported Lie algebra-function, $Q([\Xi,X])$ is a small gauge transformation.
    So let $\mathcal{H}$ be the Lie algebra of all gauge transformations, large and small, and $\mathcal{H}_{loc}$ the Lie algebra of small ones. Then the latter is an invaraint subalgebra of the former and the large algebra is the quotient $\mathcal{H}/\mathcal{H}_{loc}$ is the Lie algebra of large gauge transformations. Its elements are equivalence classes $a+\mathcal{A}_{loc}$ where $a\in\mathcal{H}$.
@@
    If $\psi$ is a gauge-noninvariant field such as a ${}^{3}$e-B or a quark field in $QCD$, and elements of $\mathcal{H}$ multiplies it, they are no longer gauge transformations , failing to fulfill requisite algebraic relations.
    Superselection Rules
    Let $\mathcal{A}_{loc}$ denote the local algebra of gauge invariant observables. Its elements commute with all elements of the gauge algebra of $\mathcal{H}$. In a customary notation , $\mathcal{H}=\mathcal{A}_{loc}^{\prime}$. So in an irreducible representation of $\mathcal{A}_{loc}$, one can diagonalise a complete commuting set of operators (CCS) from $\mathcal{H}$ and the center of its enveloping algebra. For ${}^{3}$e-B, they are picked from $Q(\Xi)$ where $\Xi$ on the boundary $S^{2}$ are maps of $S^{2}$ to $H=SU(3)_{L+R}$.
    We can illustrate these remarks by choosing the $8$ functions $\mathbb{I}\lambda_{\alpha}$ for $\Xi$ where $\mathbb{I}$ is the constant function with value $1$ on $\mathbb{R}^{3}$ and $\lambda_{\alpha}$ are the Gell-Mann matrices. Then, $Q_{\alpha}=Q(\mathbb{I}\lambda_{\alpha})$ fulfill the $SU(3)$ Lie algebra relation:
    $[Q_{\alpha},Q_{\beta}]=if_{\alpha\beta\gamma}Q_{\gamma}$ (4.15)
    for fixed Casimirs $c_{2.3}$ and a Cartan subalgebra with basis $Q_{3,8}$ , we can choose its eigenstates with eigenvalues $\tilde{\lambda}_{3,8}$ and an $SU(2)$ Casimir with eigenvalue $I(I+1)$ commuting with these $Q$’s:
    $|\tilde{\lambda}_{3},I,\tilde{\lambda}_{8},c_{2},c_{3}\rangle=|0\rangle\otimes|\tilde{\lambda}_{3},I,\tilde{\lambda}_{8},c_{2},c_{3}\rangle^{\prime}$ (4.16)
    This vector is the analogue of the vacuum vector in quantum field theory. On RHS, $|0\rangle$ is the Fock vacuum and the primed vector carries just the $SU(3)$ representation.

### v4

@@ Abstract
In 3He-B, two atoms pair in an orbital angular momentum $1$ spin triplet state above the phase transition temperature with $SO(3)\times SO(3)$ symmetry. Below the transition temperature, this symmetry is spontaneously broken to the diagonal $SO(3)$ due to spin-orbit coupling. Considerations based on effective potentials and solitons show that $SO(3)$’s get enhanced to $SU(3)$’s and the symmetry breaking is that of $G=SU(3)\times SU(3)\times U(1)$ to $H=SU(3)$. The theory of the resultant Goldtsone modes can be naturally formulated as a gauge theory of $H$. Its Gauss law is treated here and shown to lead to surface states in a container with a dynamics governed by large gauge transformations.
Observable consequences are pointed out. The transference of the analysis to the chiral model of QCD is pointed out where $SU(3)$ are the left and right chiral groups, $U(1)$ is the axial $U(1)$ , the surviving symmetry is flavour $SU(3)$ and the Goldstone modes are the pions.
I Introduction
In 3He-B , the superfluid transition takes place by a pairing of this molecule in its $P$-wave spin triplet state with another such molecule in an identical state.
The order parameter for this transition can be denoted by a complex $3\times 3$ matrix of fields $\phi=\{\phi_{sm}(x)\}$ following Mermin [1] where $SO(3)$ of spin acts on $\phi$ by left multiplication and $SO(3)$ of orbital momentum acts on $\phi$ by right multiplication. Note, the indices $s,m$ take their values in the set $\{1,2,3\}$.
@@ I Introduction
$\lim_{|\vec{x}|\rightarrow\infty}\phi(x)=\mathbb{I}$ (1)
@@
This symmetry breaking pattern is reminiscent of the chiral symmetry group $SU(3)_{L}\times SU(3)_{R}$ of massless $QCD$ with three flavours and its spontaneous breakdown to the diagonal $SU(3)_{L+R}$. As alluded to above, we can in fact bring this resemblance closer by noting that $G$-invariant potentials are traces of polynomials in $\phi^{\dagger}\phi$ of the form
$\Sigma_{n}a_{n}\mathrm{Tr}(\phi^{\dagger}\phi)^{n}.$ (2)
These potentials are invariant if $SO(3)_{L,R}$ is enlarged to $G=SU(3)_{L}\times SU(3)_{R}\times U(1)$ where $U(1)$ acts by an overall phase: $\phi\rightarrow\omega\phi,|\omega|=1$. Then the vacuum symmetry group is $H=SU(3)_{L+R}$. These are our groups of interest.
@@
$G$ does not act effectively on $\phi$. If $c=e^{\frac{2\pi i}{3}\mathbb{I}_{3}}$, with $\mathbb{I}_{3}$ being the $3\times 3$ identity matrix of $SU(3)$, generates the center of $SU(3)$, then the order $9$ discrete group $Z_{9}$ with elements diag$\{c^{m},\bar{c}^{n}\}\otimes\Omega$ with $n,m$ belonging to the set $\{1,2,3\}$ and $\Omega=e^{i(n-m)\mathbb{I}_{3}}$ acts trivially on $\phi$. Thus in quantum theory, only representations of $G$ where $Z_{9}$ becomes identity are allowed. With this proviso, we can work with $G$.
In Mermin’s review , $\phi$ is restricted to $\omega\times R$ where $R$ is real and orthogonal. But a previous work [2] has found that for discussions about solitons of ${}^{3}$e-B, it is important to consider $\phi$ to be any $3\times 3$ matrix ( with the boundary condition mentioned above). This paper will proceed in this latter direction.
II Gauge Theory of Goldstone Modes
@@
III Spontaneous Breakdown as a Gauge Theory
@@
The gauge groups $\mathcal{G},\mathcal{H}$ are the maps from spacetime to Lie groups $G,H$. In our considerations , they are the restrictions of $\mathcal{G},\mathcal{H}$ to constant time, say 00 , that will be relevant. They are denoted by the same symbols. We consider four dimensional $M$ so that the constant time slices $M_{3}$ are three dimensional. It can be $\mathbb{R}^{3}$ or the 3-ball $B_{3}$ with coordinates $x$ and the boundary $S^{2}$. The evaluation of their elements at $x$ will be denoted by $g(x)$ and $h(x)$, respectively. Often we will suppress $x$.
@@
$hT_{\alpha}h^{-1}=h_{\beta,\alpha}T_{\beta}$ (3)
@@
$v=\frac{1}{2}T(\alpha)\mathrm{Tr}(vT(\alpha)),$ (6)
@@
$A_{\mu}(g(x))=\frac{i}{2}T(\alpha)\mathrm{~{}Tr~{}}T(\alpha)g^{-1}(x)\partial_{\mu}g(x).$ (8)
@@
$A_{\mu}(gh)=h^{-1}A_{\mu}h+ih^{-1}\partial_{\mu}h.$ (10)
This follows from (5) and the identity (10) with $v_{\mu}=h^{-1}\partial_{\mu}h$.
In contrast,
$B_{\mu}(g)=\frac{i}{2}S(j)\mathrm{~{}Tr~{}}S(j)g^{-1}(x)\partial_{\mu}g(x)$ (11)
@@
$B_{\mu}(gh)=h^{-1}(x)B_{\mu}(g)h(x).$ (12)
We can now consider $\mathcal{H}$-invariant interactions like
$\mathrm{Tr~{}}F_{\mu\nu}F^{\mu\nu},\qquad\mathrm{Tr~{}}B_{\mu}B^{\mu}$ (13)
@@
$g_{L}g^{\dagger}_{R}\equiv\phi=\left\{\phi_{s,i}\right\}$ (14)
@@
IV Gauss Law and Superselection Rules
@@
$S=-\frac{1}{4g^{2}}\int d^{4}x~{}\mathrm{tr}F_{\mu\nu}F^{\mu\nu}+\int d^{4}x~{}\mathrm{tr}A^{\mu}J_{\mu}$ (15)
@@
In view of (12), $F$ transforms as $h^{-1}Fh$ under an $\mathcal{H}$ gauge transformation so that $S$ is invariant even if $h$ is time dependent.
@@
The derivation of this constraint by quantising $g$ in the action is awkward. But we can get it simply by varying $A_{0}$ with an infinitesimal time dependent gauge transformation
$\delta g=g\epsilon,~{}~{}\epsilon~{}\textrm{time dependent}$ (16)
@@
$\epsilon=\epsilon^{\alpha}(x)\,T(\alpha),$ (17)
@@
$\int d^{3}x\mathrm{tr}(A^{\mu}J_{\mu})$ (19)
to the Lagrangian which will enhance (18) by the matter term $J_{0}$.
We must check that L.H.S of (18) generates gauge transformations. and that requires the commutators of $A^{i}$ and $F_{0i}$. We can derive them as follows.
The analogue of the $pdq$ term can be read off from (15) . It comes from $\frac{1}{2g^{2}}\mathrm{tr}F_{0i}\partial_{0}A_{i}$ from varying $A_{i}$ so that as in $QCD$, $A^{i}$ and $E_{i}=\frac{1}{2g^{2}}F_{0i}$ are conjugate.
@@
$Q_{H}(X)=\int d^{3}x~{}\mathrm{tr}(D_{i}XE_{i}+XJ_{0}),$ (20)
$X=$ smooth Lie-algebra valued compactly supported functions. This is (18) smeared with test functions X.
@@
The classical Gauss law (18) has the quantum version of ‘small ’ gauge tranformations with generators $Q_{H}(X)$ which fulfill a) and b):
@@
$\left[Q_{H}(X),K\right]=0.$ (22)
Then, $X$ being compactly supported , we recover classical Gauss law by partial integration , while (22) is to ensure that observables preserve the domain defined by (21). ( There are technical issues with domains which remain to be answered.)
@@
$Q(\Xi)=\int d^{3}x~{}\mathrm{Tr}(D_{i}\Xi E_{i}+\Xi J_{0})$ (23)
where $\Xi$ are smooth functions which may not vanish at infinity ( or the boundary). These generate ‘large gauge transformations’ or the Sky group [7]. As shown in previous papers, $Q(\Xi)$ commute with local observables due to local commutation relations , but in non-abelian theories such as ${}^{3}$e-$B$ or $QCD$, they may not mutually commute. This has specific consequences for surface states as we shall see.
@@
    $\left[Q(\Xi),Q(\Xi^{\prime})\right]=iQ([\Xi,\Xi^{\prime}]).$ (24)
@@
    Adding a small gauge transformation $Q_{H}(X)$ to $Q(\Xi)$ is still a large gauge transformation $Q(\Xi+X)$ which has the same action on a quantum state in view of (21):
    $Q(\Xi+X)|\cdot\rangle=\bigg{(}Q(\Xi)+Q(X)\bigg{)}|\cdot\rangle=Q(\Xi)|\cdot\rangle$ (25)
Further since $[\Xi,X]$ is a compactly supported Lie algebra-function, $Q([\Xi,X])$ is a small gauge transformation. ( As $X$ has a compact support, so does $[\Xi,X]$.)
So, let ${\underline{\mathcal{H}}}$ be the Lie algebra of all gauge transformations, large and small, and $\underline{\mathcal{H}_{loc}}$ the Lie algebra of small ones. Then the latter is an invaraint subalgebra of the former and the large algebra is the quotient ¯$\underline{\mathcal{H}}/\underline{\mathcal{H}_{loc}}$. Its elements are equivalence classes $a+\underline{\mathcal{H}_{loc}}$ where $a\in\underline{\mathcal{H}}$.
@@
If $\psi$ is a gauge-noninvariant field such as a ${}^{3}$e-B or a quark field of $QCD$, and elements of $\underline{\mathcal{H}}$ multiplies it, they are no longer gauge transformations , failing to fulfill requisite algebraic relations.
IV.1 Superselection Rules
Let $\mathcal{A}_{loc}$ denote the local algebra of gauge invariant observables. Its elements commute with all elements of the gauge algebra $\underline{\mathcal{H}}$.
In a customary notation , ¯$\underline{\mathcal{H}}=\mathcal{A}_{loc}^{\prime}$. So in an irreducible representation of $\mathcal{A}_{loc}$, one can diagonalise a complete commuting set of operators (CCS) from $\underline{\mathcal{H}}$ and the center of its enveloping algebra. For ${}^{3}$e-B confined in a ball, they are picked from $Q(\Xi)$ where $\Xi$ on the boundary $S^{2}$ of $\mathbb{R}^{3}$ or a spherical container of ${}^{3}$e-B are maps of $S^{2}$ to $H=SU(3)_{L+R}$.
We can illustrate these remarks by choosing the $8$ functions $\mathbb{I}\lambda_{\alpha}$ for $\Xi$ where $\mathbb{I}$ is the constant function with value $1$ on $\mathbb{R}^{3}$ or the ball and $\lambda_{\alpha}$ are the Gell-Mann matrices. Then, $Q_{\alpha}=Q(\mathbb{I}\lambda_{\alpha})$ fulfill the $SU(3)$ Lie algebra relation:
$[Q_{\alpha},Q_{\beta}]=if_{\alpha\beta\gamma}Q_{\gamma}.$ (26)
We can diagonalise $Q_{3,8}$ and $Q_{1}^{2}+Q_{2}^{2}+Q_{3}^{2}(=I(I+1))$ after fixing Casimirs to values $c_{2},c_{3}$ to obtain the vector
$|\tilde{\lambda}_{3},I,\tilde{\lambda}_{8},c_{2},c_{3}\rangle=|0\rangle\otimes|\tilde{\lambda}_{3},I,\tilde{\lambda}_{8},c_{2},c_{3}\rangle^{\prime}.$ (27)
This vector is the analogue of the vacuum vector in quantum field theory. On RHS, $|0\rangle$ is the Fock vacuum or the ground state and the primed vector carries just the $SU(3)$ representation.